In the preparation of alkanes by the reduction of alkyl halides, we observe the following reaction:


R-X + 2H --------> R-H + X2


X=Halogen, R=Alkyl group


Can you explain the mechanism of this reaction?

Grade:11

1 Answers

vikas askiitian expert
13 years ago

R-X + 2H --------------> R-H + X2

R-X , H-H bond is cleaved homolatically by using light (high temperature) to form free radicals...

R-X = R. + X.

H2 = H. + H.

 R.+R. = R-R

 X. + X. = X2

 R.+H. = R-H

this reaction is free radical substitution ...